# Fuzzy to Clear: Elucidating the Threat Hunter Cognitive Process and Cognitive Support Needs

### v2

@@ 1 Introduction
The complexity and challenges of the cybersecurity landscape have dramatically increased over the past several years. In 2023 alone, 2,365 cyber attacks were reported in the United States, with an estimated 343,338,964 victims [^1] https://www.forbes.com/advisor/education/it-and-tech/cybersecurity-statistics/, while another global threat report indicates the fastest observed breakout time for interactive eCrime intrusion was only 2 minutes and 7 seconds [^2] https://www.crowdstrike.com/en-us/global-threat-report/. As the frequency, speed, and severity of cyber threats continue to rise, it is more crucial than ever for organizations to implement robust security measures. One such measure is the practice of threat hunting, which plays a critical role in combating increasingly sophisticated and evolving cyber threats [1].
@@
Threat hunting is the proactive process of searching IT infrastructure for signs of malicious activity and suspicious behaviours that have evaded existing security defences [2]. It involves comprehensive data analysis from various sources, including network traffic, system logs, endpoint telemetry, and threat intelligence feeds, to uncover potential threats [1]. Threat hunting can be performed by different cybersecurity roles, such as security analysts. However, a specialized full-time role has emerged: the threat hunter. Throughout our paper, when we say threat hunter (TH), we refer to the subset of cybersecurity professionals for whom threat hunting is their primary role.
@@
More broadly, multiple studies describe the challenges faced by these highly skilled cybersecurity professionals, who are facing growing demand for their time and prone to burnout [3, 4]. Examples of challenges include the cognitive workload required to sift through large amounts of data [5] and the need to develop situational awareness of complex systems [6]. Additionally, these professionals must navigate these challenges during the already difficulty process of collaborating with their peers, clients and/or stakeholders [7]. Unlike reactive security strategies that rely on automated threat detection systems’ alerts to act, THs need to proactively examine extensive and often complex logs and user activities to identify stealthy threats that may be undetectable by conventional methods [8]. This complicates the TH role beyond the challenges outlined above (as applicable to all cybersecurity professionals). These challenges require further investigation into threat hunting practices and the development of effective support for human-in-the-loop approaches throughout this process [8, 1].
@@
In this complex context, designing strategies and developing innovative tools to support THs becomes crucial. The problem is that designers lack studies that capture THs’ cognitive processes or challenges that could aid the design process. Although we can find multiple tools, methodologies and models supporting the threat hunting process [2, 8, 9, 1], there is still a lack of studies and empirical evidence that captures the THs’ cognitive processes and the challenges they struggle with (similar limitations are also discussed by other authors on human-centered cybersecurity [10, 7]).
@@ 2.1 Threat Hunting
During structured hunting, generating high-quality hypotheses is paramount [2], and is often a manual process requiring knowledge and expertise [8]. However, this is not always the case; unstructured hunting, for example, does not involve the creation of hypotheses [9, 2]. Threat hunters may also belong to broader security teams or larger organizations [14, 2]. In general, these cybersecurity teams require skills such as an ability to communicate their findings to superiors and shared situational awareness [7]. Situational awareness is the concept of “knowing what is going on around you” [15]. Much of the research into situational awareness in cybersecurity has a focus on using AI and ML to automate the tasks of security analysts, rather than focusing on supporting the needs of the humans doing the job [16, 7, 6]. Additional skills are required by threat hunting teams in particular, for which Maxam and Davis [9] provide a detailed account.
Automation helps THs to be more efficient, however, threat actors are bypassing automated systems and still infiltrating organizations. Gutzwiller et al. [16] acknowledges that human judgment remains crucial for validating and making complex decisions. Automated processes still need to be supervised by humans and adjusted based on human insights, emphasizing the importance of a human-in-the-loop approach. This claim is also supported by recent studies in the threat hunting scope such as Nour et al. [8] and Mahboubi et al. [1].
Despite significant advancements in the field, THs continue to face numerous obstacles that hinder the effective detection and prevention of malicious activities. These challenges highlight the complexity and ever-evolving nature of cyber threats, as well as the gaps in current methodologies, technologies, and analyst skills [1]. In a recent systematic literature review, Mahboubi et al. [1] summarize key difficulties in threat hunting, including: (a) a lack of labeled data, (b) imbalanced datasets, (c) multiple sources of log data, (d) adversarial techniques, and (e) a shortage of human experts and data intelligence. Understanding these challenges is essential to improving the effectiveness and efficiency of threat hunting practices.
Given the existing literature and the challenges THs face, there is a gap in the knowledge surrounding the behaviors and cognitive processes of THs, and how to address the related challenges. It is crucial to understand how THs build and utilize mental models during threat hunting sessions in order to address this gap. With our research, we aim to contribute an understanding of how THs build and utilize mental models during threat hunting sessions, and propose five actionable design propositions that can inform the design of tools to better support THs.
@@ 2.2 Models and their Applications in Security
In this subsection, we overview various process models used in cybersecurity. For simplicity, we consider everything we discuss to be a process model, where we use Curtis et al. [17]’s definition as “an abstract description of an actual or proposed process that represents selected process elements that are considered important to the purpose of the model and can be enacted by a human or machine”.
@@ 2.2.1 Models of Threat Actor Behaviors
Models are used in cybersecurity to make frequently-used attack patterns and vectors more accessible. Three of the most popular threat actor behavior models are the MITRE ATT&CK [18], Pyramid of Pain [19], and Cyber Kill Chain [20], which all describe common techniques and behaviors of cyber attackers. The process of a threat hunting team will vary from team to team depending on the different hunting approaches: structured hunting (hypothesis-based) vs. unstructured hunting (data-driven) [9, 2], and the structure of the organization: government vs. private sector [9].
@@ 2.2.2 Workflow Models for Threat Hunting
Three notable workflow models from Gunter and Seitz [21], van Os et al. [2], and Maxam and Davis [9] describe the processes used by THs to conduct threat hunting tasks. Gunter and Seitz propose a six stage workflow model for the threat hunting process that defines a “proactive, analyst-driven process to search for attacker [tactics, techniques, and procedures (TTP)] within an environment” [21]. Built on previous work like the Cyber Kill Chain [20], the authors describe purpose, scope, equip, plan review, execute, and feedback as the six necessary steps for completing a cycle in the threat hunting process [21]. The model is structured as a cycle to “ensure that the results and lessons learned from previous hunts influence future hunts.” [21]. The model ensures that THs have a standardized methodology for hunting and maintain a “a closer focus on” [21] their objectives.
Van Os et al. propose the TaHiTI (Targeted Hunting integrating Threat Intelligence) workflow model that provides “a focused and risk-driven” [2] methodology, and models a hypothesis-based method of threat hunting [2]. Three phases, initiate, hunt, and finalize, are described in the model [2]. First, a trigger in the initiation phase results in the creation of the hunting investigation abstract (often attacker behavior is identified through MITRE ATT&CK) [2]. Secondly, in the hunting phase, THs complete two activities: define / refine: clarifying details for the hunt and execute: retrieving and analyzing data [2]. In the third finalization phase, results from the previous step are processed and the hunt’s findings are documented [2].
Maxam and Davis describe the threat hunting process of government security teams motivated by other workflow models like TaHiTi, because TaHiTI did not capture threat hunting processes described by Maxam and Davis’ participants [9]. Maxam’s model splits the threat hunting process into seven stages: begin hunt, mission planning, collect intelligence, pre-mission activities, manual and automatic analysis loops, and end mission [9].
These three models capture various perspectives of the workflows of THs. What they do not capture are the steps THs take to build, update, and share their mental models during hunting. This is the gap our study will address.
@@ 2.2.3 Models For Cognitive Tasks
Other models in cybersecurity focus on cognitive aspects, representing how analysts’ minds process information while perceiving, comprehending, and responding to threats. In fields such as cognitive security [22] and Cyber Situation Awareness [23], these cognitive task models form the basis of proposed strategies to automate the threat hunting process undertaken by cybersecurity analysts [10, 24, 25, 26]. Damico et al. previously used a cognitive task analysis approach to construct a generalized workflow of how information assurance analysts build situational awareness and respond to threats [27].
@@ 2.2.4 Models From Other Fields
Cybersecurity has employed techniques based on popular models originating from other fields, such as the OODA (Observe, Orient, Decide, Act) originally developed by John Boyd [28] to enhance situational awareness in military aviation operations. The OODA loop is an iterative and adaptive process, crucial for maintaining a competitive edge in dynamic environments such as the cyber threat landscape. OODA is applicable to the threat hunting process where rapid decision-making and continuous adaptation are essential.
Models from other fields such as in programming [29, 30, 31, 32, 33], engineering education [34], social science [35], Human-Computer Interaction (HCI) [36], medical education [37], science education [38], and business [39] can also support the development of better understanding, performance, and decision-making in threat hunting. For example, Heinonen et al.[31] and LaToza et al.[32] explore how developers maintain mental models of code, emphasizing communication and implicit knowledge similar to how THs need to share information with their teammates. In medical education, cognitive load theory is used to reduce cognitive load [37], a challenge also faced by THs. Science education employs mental models to create more intuitive and effective tools [38]. In business, models are used and shared within organizations to improve shared cognition and overall performance [39]. These related studies highlight the importance of using models for improving understanding and problem-solving and inspiring the creation of models that support threat hunting.
Despite the extent of the related literature, existing models fail to explain how THs build, use, and share their mental models. This gap demonstrates the need for deeper research into the behavioral, cognitive, and collaborative aspects of threat hunting [40, 16, 12]. Our study aims to extend existing research by developing a model that describes the process of building a mental model during a threat hunting session. Building such a model involves understanding the key cognitive tasks, their sequence, and how they contribute to the formation and utilization of mental models in dynamic and complex environments. In the next subsection, we look at related work on mental models in more detail as we explore the construction of TH mental models in our work.
@@ 2.3 Mental Models
Mental models are defined by Human-Computer Interaction (HCI) researcher Don Norman as the conceptual models residing in people’s minds; they are (typically highly simplified) explanations for how things in the world work [41]. Mental models form through interactions with the environment, others, and technology [42]. Mental models are also characterized as evolving over time, often containing inaccuracies, and are limited by the user’s technical background or previous experiences [42]. Two people will not necessarily share the same mental model, and one person may hold multiple mental models of a single item to represent its different functions [41].
Mental models are well-adopted across many fields such as medical education, science education, and business [37, 38, 39], including HCI [36]. Mental models have been studied in the cybersecurity context [43, 44] as well. The existing studies of cybersecurity mental models address how end users perceive threats [44], help cyber analysts conceptualize threat actors and threats themselves [27], and enable automation of TH responsibilities. However, none of the cybersecurity research of mental models address how THs build a mental model while hunting.
Mahboubi et al. [1] explore through a systematic review the role of adaptive cognitive processes in threat hunting workflows, highlighting how analysts combine hypothesis-driven approaches with iterative refinement to address dynamic and evolving threats. Their findings underscore the importance of mental models in guiding decision-making and aligning human insights with AI-driven tools, emphasizing a human-centered approach to navigating complex cybersecurity environments.
Fortuna et al. [45] also discusses how cognitive models in cybersecurity can be leveraged to support professionals in high-stakes environments. This aligns with the need for tools that account for the cognitive processes of threat hunters, helping them form and adapt mental models to make informed decisions during the hunting process. Cyberpsychology, as discussed by Norman [46], highlights the interaction between human cognition and technology, underscoring the importance of designing tools that align with users’ cognitive processes. This human-centered approach is essential for improving usability, reducing cognitive load, and enhancing situational awareness for professionals working in cybersecurity environments, similarly to the gap in situational awareness for cybersecurity noted by Gutzwiller [16].
Murimi et al. note that in order for mental models to be effective in the context of cybersecurity they require “a human-centered foundation, with consideration of the technology, situational awareness, and human behavior” [44]. In HCI, mental models are pivotal for enhancing system design and usability. Mental models help designers create interfaces that align with user expectations, thereby reducing cognitive load and increasing efficiency [36]. By understanding these mental models, designers can predict user interactions and design more intuitive and user-friendly interfaces that not only improve usability, but also enhance accessibility. Understanding mental models and how they are constructed is crucial for the development of tools that support their users. Our model of the threat hunter process of building a mental model and proposed design propositions support the development of tools that better support THs.
@@ 3.1 Study Process and Overview
The study protocol was reviewed and approved by the Human Research Ethics Office at the University of Victoria (project application number 21-0601).
@@ 3.1.1 Industry Partner
This study was conducted as part of a research project in collaboration with OpenText [^3] https://www.opentext.com/products/cyber-security. Our industry partner is a global market leader for information management software that offers cybersecurity consulting and technology services. One of the products offered by OpenText is ArcSight Intelligence, described as an advanced threat-detection tool that uses user entity behaviour analytics (UEBA) and unsupervised machine learning models to detect behavioural anomalies across the organization [^4] https://www.opentext.com/products/arcsight-intelligence.
@@ 3.1.3 Observation procedure and tools
The observation sessions encompass all six formal meetings the research team had with the four participants (see Fig. 1-b). The participants were encouraged to conduct their threat hunting activities while maintaining their natural environment as much as possible. However, in most sessions, the participants chose to share their process based on previous investigations they performed rather than the task planned for their current shift. This approach of replaying a threat hunting session (rather than real-time hunting) allowed them to share more diverse cases, but is also an approach they use while training new members to their team (so as such is quite representative of the typical process but may be more constrained to a certain time frame).
@@
We attended the observation sessions without predefined hypotheses, allowing us to take notes freely without following a specific template or set of guiding questions (i.e., an exploratory mindset of gathering an understanding of the threat hunting practices and their related cognitive processes). Additionally, we decided to limit our questions (or any interruptions for clarification) to only a few to avoid disruptions to the activities and natural flow of the work being conducted by the participants. Thus, most of the questions were noted to be clarified later (during validation and member checking sessions, explained in Subsection 3.3).

### v3

@@ 1 Introduction
The complexity and challenges of the cybersecurity landscape have dramatically increased over the past several years. In 2023 alone, 2,365 cyber attacks were reported in the United States, with an estimated 343,338,964 victims, [^1] https://www.forbes.com/advisor/education/it-and-tech/cybersecurity-statistics/ while another global threat report indicates the fastest observed breakout time for interactive eCrime intrusion was only 2 minutes and 7 seconds. [^2] https://www.crowdstrike.com/en-us/global-threat-report/ As the frequency, speed, and severity of cyber threats continue to rise, it is more crucial than ever for organizations to implement robust security measures. One such measure is the practice of threat hunting, which plays a critical role in combating increasingly sophisticated and evolving cyber threats [1].
@@
Threat hunting is the proactive process of searching IT infrastructure for signs of malicious activity and suspicious behaviours that have evaded existing security defences [2]. It involves comprehensive data analysis from various sources, including network traffic, system logs, endpoint telemetry, and threat intelligence feeds, to uncover potential threats [1]. Threat hunting can be performed by different cybersecurity roles, such as security analysts, who may be part of dedicated teams or have other organizational responsibilities [3]. However, a specialized full-time role has emerged: the threat hunter. Throughout our paper, when we say threat hunter (TH), we refer to the subset of cybersecurity professionals for whom threat hunting is their primary role.
@@
More broadly, multiple studies describe the challenges faced by these highly skilled cybersecurity professionals, who are facing growing demand for their time and prone to burnout [4, 5]. Examples of challenges include the cognitive workload required to sift through large amounts of data [6] and the need to develop situational awareness of complex systems [7]. Additionally, these professionals must navigate these challenges during the already difficulty process of collaborating with their peers, clients and/or stakeholders [8]. Unlike reactive security strategies that rely on automated threat detection systems’ alerts to act, THs need to proactively examine extensive and often complex logs and user activities to identify stealthy threats that may be undetectable by conventional methods [9]. This complicates the TH role beyond the challenges outlined above (as applicable to all cybersecurity professionals). These challenges require further investigation into threat hunting practices and the development of effective support for human-in-the-loop approaches throughout this process [9, 1, 3].
@@
In this complex context, designing strategies and developing innovative tools to support THs becomes crucial. The problem is that designers lack studies that capture THs’ cognitive processes or challenges that could aid the design process. Although we can find multiple tools, methodologies and models supporting the threat hunting process [2, 9, 10, 1, 3], there is still a lack of studies and empirical evidence that captures the THs’ cognitive processes and the challenges they struggle with (similar limitations are also discussed by other authors on human-centered cybersecurity [11, 8]).
@@ 2.1 Threat Hunting
During structured hunting, generating high-quality hypotheses is paramount [2], and is often a manual process requiring knowledge and expertise [9]. However, this is not always the case; unstructured hunting, for example, does not involve the creation of hypotheses [10, 2] (we discuss threat hunting processes and workflows further in Subsection 2.2.2).
Threat hunters may also belong to broader security teams or larger organizations [15, 2]. In general, these cybersecurity teams require skills such as an ability to communicate their findings to superiors and shared situational awareness [8]. Situational awareness is the concept of “knowing what is going on around you” [16]. Much of the research into situational awareness in cybersecurity has a focus on using AI and ML to automate the tasks of security analysts, rather than focusing on supporting the needs of the humans doing the job [17, 8, 7]. Additional skills are required by threat hunting teams in particular, for which Hill et al. [18], Maxam and Davis [10], and Badva et al. [3] provide a detailed account—three recent studies conducting interviews in the scope of threat hunting.
Automation plays a significant role in supporting threat hunting activities, and recent advancements—such as the automated generation of attack hypotheses [19, 20]—have contributed meaningfully to this domain. However, automation alone cannot address all the complexities of threat hunting; human expertise remains essential, as emphasized by different authors such as [9, 1, 3]. This underscores the importance of a human-in-the-loop approach. Still, despite this importance, many automated tools and related studies often overlook the impact of automation on human interaction and the cognitive dimensions of threat investigation—a gap also highlighted by [3].
THs face numerous obstacles that hinder the effective detection and prevention of malicious activities. In a recent systematic literature review, Mahboubi et al. [1] summarize key difficulties in threat hunting, including: (a) a lack of labeled data, (b) imbalanced datasets, (c) multiple sources of log data, (d) adversarial techniques, and (e) a shortage of human experts and data intelligence. These challenges highlight the complexity and ever-evolving nature of cyber threats, as well as the gaps in current methodologies, technologies, and analyst skills, as also noted by other authors [3, 10, 18, 9]. Understanding these challenges is essential to improving the effectiveness and efficiency of threat hunting practices. In particular, there is a significant opportunity for further research and tool development focused on the behaviors and cognitive processes of threat hunters, which remain underexplored.
@@ 2.2 Models and their Applications in Security
In this subsection, we overview various process models used in cybersecurity. For simplicity, we consider everything we discuss to be a process model, where we use Curtis et al. [21]’s definition as (p. 76) “an abstract description of an actual or proposed process that represents selected process elements that are considered important to the purpose of the model and can be enacted by a human or machine”.
@@ 2.2.1 Models of Threat Actor Behaviors
Models are used in cybersecurity to make frequently-used attack patterns and vectors more accessible. Three of the most popular threat actor behavior models are the MITRE ATT&CK [22, 23], Pyramid of Pain [24], and Cyber Kill Chain [25], which all describe common techniques and behaviors of cyber attackers. The process of a threat hunting team will vary from team to team depending on the different hunting approaches: structured hunting (hypothesis-based) vs. unstructured hunting (data-driven) [10, 2], and the structure of the organization: government vs. private sector [10].
@@ 2.2.2 Workflow Models for Threat Hunting
One of the earliest references to structured processes or workflow models in threat hunting are the works of Gunter and Seitz [26] and van Os et al.[2]. Building on foundational concepts such as the Cyber Kill Chain[25], Gunter and Seitz [26] propose a six-step cyclical model including purpose, scope, equip, plan review, execute, and feedback. Their model [26] emphasizes continuous improvement by ensuring that insights from previous hunts inform future ones. While it aims to provide threat hunters with a standardized methodology and a stronger focus on objectives, the model remains relatively abstract.
In contrast, van Os et al. [2] offer a more detailed and operationally rich approach through the TaHiTI (Targeted Hunting integrating Threat Intelligence) workflow model. TaHiTI follows a hypothesis-driven, risk-focused methodology and defines three distinct phases. First, in the Initiate phase, a trigger (e.g., a security incident, domain knowledge, or input from the MITRE ATT&CK framework) leads to the creation of a hunting investigation abstract, including an initial hypothesis and priority. Next, the Hunt phase involves two key activities: define/refine, where the hypothesis is clarified, and execute, where data is gathered and analyzed. The Finalize phase involves processing results, documenting findings, and handing them off to relevant processes (e.g., incident response, threat intelligence generation, or vulnerability management).
Empirical studies offering interview-based insights into the threat hunting process have only recently emerged (e.g., [10, 3]), introducing a descriptive process instead of a prescriptive approach as in the first two mentioned earlier. Maxam and Davis[10] outline the process across seven distinct stages: begin hunt, mission planning, collect intelligence, pre-mission activities, manual and automatic analysis loops, and end mission, providing detailed contextual information for each phase. Conversely, Badva et al. [3] present a higher-level conceptual model structured as a cyclical workflow. Their workflow model [3] begins with the selection of a Threat Hunting Method, which includes approaches as use-case-based, predefined scenarios or patterns of suspicious activities to identify and investigate known threats and attack patterns, intel-based, leveraging technical threat intelligence, or random-based hunting, without a predefined plan. This phase is followed by phases of Pre-Hunt Planning, Data Collection & Preparation, Hunting & Validation, and Remediation & Reporting, before looping back to the planning phase.
Together, these models ([26, 2, 10, 3]) offer diverse perspectives on the workflows followed by threat hunters. However, they do not address the underlying cognitive processes—specifically, how threat hunters build, refine, and share their mental models throughout the hunting process.
@@ 2.2.3 Models for Cognitive Tasks
Other models in cybersecurity focus on cognitive aspects, representing how analysts’ minds process information while perceiving, comprehending, and responding to threats. In fields such as cognitive security [27] and cyber situation awareness [28], these cognitive task models form the basis of proposed strategies to automate the threat hunting process undertaken by cybersecurity analysts [11, 29, 30, 31]. D’Amico et al. [32] previously used a cognitive task analysis approach to construct a generalized workflow of how information assurance analysts build situational awareness and respond to threats.
@@ 2.2.4 Models from Other Fields
Cybersecurity has employed techniques based on popular models originating from other fields, such as the OODA (Observe, Orient, Decide, Act) originally developed by John Boyd [33] to enhance situational awareness in military aviation operations. The OODA loop is an iterative and adaptive process, crucial for maintaining a competitive edge in dynamic environments such as the cyber threat landscape. OODA’s concepts apply to defending from cyberthreats [34], and therefore can also be used in the threat hunting process.
Models from other fields such as in programming [35, 36, 37], Human-Computer Interaction (HCI) [38], medical education [39], science education [40], and business [41] can also support the development of better understanding, performance, and decision-making in threat hunting. For example, Heinonen et al.[35] and LaToza et al.[36] explore how developers maintain mental models of code, emphasizing communication and implicit knowledge similar to how THs need to share information with their teammates. In medical education, cognitive load theory is used to reduce cognitive load [39], a challenge also faced by THs. Science education employs mental models to create more intuitive and effective tools [40]. In business, models are used and shared within organizations to improve shared cognition and overall performance [41]. These related studies highlight the importance of using models for improving understanding and problem-solving and inspiring the creation of models that support threat hunting.
In the next subsection, we look at related work on mental models in more detail as we explore the construction of TH mental models in our work.
@@ 2.3 Mental Models
Mental models are defined by HCI researcher Don Norman as the conceptual models residing in people’s minds; they are (typically highly simplified) explanations for how things in the world work [42]. Mental models form through interactions with the environment, others, and technology [43]. Mental models are also characterized as evolving over time, often containing inaccuracies, and are limited by the user’s technical background or previous experiences [43]. Two people will not necessarily share the same mental model, and one person may hold multiple mental models of a single item to represent its different functions [42].
In the cybersecurity context, existing studies of mental models address how end users perceive threats [44, 45], help cyber analysts conceptualize threat actors and threats themselves [32], and enable automation of TH responsibilities. Mahboubi et al.[1] highlight how adaptive cognitive processes, including iterative refinement and hypothesis-driven approaches, help analysts address evolving threats, underscoring the importance of mental models in guiding decision-making and integrating human insights with AI tools.
Fortuna et al.[46] and Norman [47] similarly emphasize the role of cognitive models and human-technology interaction in supporting professionals in complex environments, with a human-centered approach essential for reducing cognitive load and enhancing situational awareness, as also noted by Gutzwiller et al.[17]. Murimi et al.[44] stress that effective cybersecurity mental models require a human-centered foundation that accounts for technology, situational awareness, and human behavior. This aligns with insights from HCI, where, as Hu et al. [38] note, mental models help design more intuitive and accessible systems. Understanding how these models are constructed is crucial for developing tools that truly support their users.
Despite the breadth of related literature, including studies on adaptive cognitive processes [1], human-technology interaction [46, 47], and the need for human-centered foundations in cybersecurity mental models [44, 38], existing research still fails to explain how THs specifically build, use, and share their mental models during active hunting sessions. This gap underscores the need for deeper investigation into the behavioral, cognitive, and collaborative aspects of threat hunting. Addressing this gap, our study focuses on the key cognitive tasks and sequences involved in constructing mental models to navigate complex and dynamic cybersecurity environments.
@@ 3.1 Study Process and Overview
The study protocol was reviewed and approved by the Human Research Ethics Office at the University of Victoria, Canada (project application number 21-0601).
@@ 3.1.1 Industry Partner
This study was conducted as part of a research project in collaboration with OpenText. [^3] https://www.opentext.com/products/cyber-security Our industry partner is a global market leader in information management software offering a variety of cybersecurity products and services. One of the products offered by OpenText is ArcSight Intelligence, described as an advanced threat-detection tool that uses user entity behaviour analytics (UEBA) and unsupervised machine learning models to detect behavioural anomalies across the organization. [^4] https://www.opentext.com/products/arcsight-intelligence
@@ 3.1.3 Observation procedure and tools
The observation sessions encompass all six formal meetings the research team had with the four participants (see Fig. 1-b). The participants were encouraged to conduct their threat hunting activities while maintaining their natural environment as much as possible. However, in most sessions, the participants chose to share their process based on previous investigations they performed rather than the task planned for their current shift. This retrospective approach—replaying a past threat hunting session rather than engaging in real-time hunting—allowed them to present a wider range of cases. These cases included hunting sessions triggered by different reasons and hunting approaches (such as use-case or intel-based, see Subsection 2.2.2 for more details), and not only successful threat discoveries but also their broader investigative process, including marking events that ultimately did not lead to real threats or attacks. This approach reflects how they typically train new team members, making it a representative illustration of their standard practices (though potentially time-constrained).
@@
We attended the observation sessions without predefined hypotheses, allowing us to take notes freely without following a specific template or set of guiding questions (i.e., an exploratory mindset of gathering an understanding of the threat hunting practices and their related cognitive processes). Additionally, we decided to limit our questions or any interruptions for clarification to only a few to avoid disruptions to the activities and natural flow of the work being conducted by the participants (i.e., participants were not answering predefined questions as would be the case in an interview without the observations). Thus, most of the questions were noted to be clarified later (during validation and member checking sessions, explained in Subsection 3.3).